Quarter Gallons of Chemical Solution
1 550
2 490
3 510
4 590
5 595
6 550
7 565
8 610
9 635
10 565
11 572
12 655

Using the data conduct the following:

a. Starting with a forecast of 550 for quarter 1, forecast the demand for quarter 13 using exponential smoothing with α = 0.6. Plot the actual data and the exponentially smoothed forecasts on the same graph. SHOW FORMULA

b. Forecast demand from quarter 2 to quarter 13 using the naïve approach. Plot your forecasts on the graph developed in part a. SHOW FORMULA

c. Looking at the graph in part b, which forecasting method seems to be more appropriate?

d. Determine the MAD (disregard quarter 1 from computations) to confirm your answer in part c. SHOW FORMULA IF USED.
Answer #1

EXPONENTIAL SMOOTHING

F(T) = F(T-1) + (ALPHA * (A(T-1) - F(T-1)))

FORECAST 2 = 550 + (0.6 * (550 - 550)) = 550

FORECAST 3 = 550 + (0.6 * (490 - 550)) = 514

FORECAST 4 = 514 + (0.6 * (510 - 514)) = 512

FORECAST 5 = 512 + (0.6 * (590 - 512)) = 559

FORECAST 6 = 559 + (0.6 * (595 - 559)) = 581

FORECAST 7 = 581 + (0.6 * (550 - 581)) = 562

FORECAST 8 = 562 + (0.6 * (565 - 562)) = 564

FORECAST 9 = 564 + (0.6 * (610 - 564)) = 592

FORECAST 10 = 592 + (0.6 * (635 - 592)) = 618

FORECAST 11 = 618 + (0.6 * (565 - 618)) = 586

FORECAST 12 = 586 + (0.6 * (572 - 586)) = 578

FORECAST 13 = 578 + (0.6 * (655 - 578)) = 624


MAD


PERIOD

ACTUAL DEMAND

FORECAST

DEVIATIONS

ABSOLUTE DEVIATION

1

550

2

490

550

-60

60

3

510

514

-4

4

4

590

512

78

78

5

595

559

36

36

6

550

581

-31

31

7

565

562

3

3

8

610

564

46

46

9

635

592

43

43

10

565

618

-53

53

11

572

586

-14

14

12

655

578

77

77

SIGMA

445


MAD = SIGMA(ABS DEV) / N

MAD = 445 / 11 = 40.45


NAIVE FORECAST

FORECAST = DEMAND FOR PREVIOUS PERIOD


PERIOD

DEMAND

NAIVE FORECAST

1

550

2

490

550

3

510

490

4

590

510

5

595

590

6

550

595

7

565

550

8

610

565

9

635

610

10

565

635

11

572

565

12

655

572

13

655



MAD


PERIOD

ACTUAL DEMAND

FORECAST

DEVIATIONS

ABSOLUTE DEVIATION

1

550

2

490

550

-60

60

3

510

490

20

20

4

590

510

80

80

5

595

590

5

5

6

550

595

-45

45

7

565

550

15

15

8

610

565

45

45

9

635

610

25

25

10

565

635

-70

70

11

572

565

7

7

12

655

572

83

83

SIGMA

455



MAD = 455 / 11 = 41.36



BASED ON THE MAD, THE EXPONENTIAL SMOOTHING FORECAST IS BETTER.